​FarmCo, Inc. follows a policy of paying out cash dividends equal to the residual amount that remains after funding 40 percent of its planned capital expenditures. The firm tries to maintain a 40 percent debt and 60 percent equity capital structure and does not plan on issuing more stock in the coming year.​ FarmCo's CFO has estimated that the firm will earn $ 19 million in the current year.

a. If the firm maintains its target financing mix and does not issue any equity next​ year, what is the most it could spend on capital expenditures next year given its earnings​ estimate?

b. If​ FarmCo's capital budget for next year is $ 17 ​million, how much will the firm pay in dividends and what is the resulting dividend payout​ percentage?
Answer #1
a) The maximum it could spend in capital expenditure = 19/60% = $          31.67 million
b) Portion of capital budget to be met out of retaied earnings = 17*40% = $             6.80 million
Balance of NI available for dividends = 19-6.8 = $          12.20 million
Resulting dividend payout percentage = 12.20/19.00 = 64.21%